1) Within February 2011: Test the FEE on P3S2 and P4S2 (Chambers which are on the floor of the cave (03.02.2011). Next week, on Monday (06.02) Torsten will prepare the ground cable and the POF connector with 2-4 meters POF cables. Attilio will get data and test the macros for the data analysis.

2) 06.02-11.02 Prepare standalone test in the electronics lab, in order to check the damaged MB and OEP. If it is possible one should recover them.

3) Within March 2011: Electronics test of all sectors of P2, P3, P4 and compare with the list with Vladimir and Olga provided (Nov10 beam time). Understand the problems of the MB and OEP.

4) when??: Electronics test of all sectors of P1.

5) when??: Test the FEE which is in the lab: 6 long MBs, 6 short MBs and try to recover the them.

6) when??: Test the chamber with pulser and ask Vladimir to cross check.

7) when??: Implement rotating mask for the calibration. Test in lab and in the cave.

8) 06.02-11.02: Find a solution to glue the OEP to the MB. Check all voltages on the FANPV and label the voltages which had been adjusted.

9) when??: Check the all OEPBs are properly connected. Check all low voltage connectors (in particular on the FANPW side) and substitute the broken one or glue them?!

-- AttilioTarantola - 03 Feb 2011
Topic revision: r7 - 2011-02-11, AttilioTarantola
Copyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki Send feedback | Imprint | Privacy Policy (in German)